# LoVR: A Benchmark for Long Video Retrieval in Multimodal Contexts

### v1

@@
Abstract
Long videos contain a vast amount of information, making video-text retrieval an essential and challenging task in multimodal learning. However, existing benchmarks suffer from limited video duration, low-quality captions, and coarse annotation granularity, which hinder the evaluation of advanced video-text retrieval methods. To address these limitations, we introduce LoVR, a benchmark specifically designed for long video-text retrieval. LoVR contains 467 long videos and over 40,804 fine-grained clips with high-quality captions. To overcome the issue of poor machine-generated annotations, we propose an efficient caption generation framework that integrates VLM automatic generation, caption quality scoring, and dynamic refinement. This pipeline improves annotation accuracy while maintaining scalability. Furthermore, we introduce a semantic fusion method to generate coherent full-video captions without losing important contextual information. Our benchmark introduces longer videos, more detailed captions, and a larger-scale dataset, presenting new challenges for video understanding and retrieval. Extensive experiments on various advanced embedding models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research. We release the code and dataset link at https://github.com/TechNomad-ds/LoVR-benchmark
Figure 1: Example of video data from LoVR. A long video consists of multiple clips, three of which are shown on the left. The corresponding caption is displayed on the right, describing each clip’s scene and atmosphere, their inter-linkage, and an overall theme summary.
Table 1: Comparison of statistics of video-text benchmarks.
Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator Theme Info.
MSR-VTT xu2016msr 1,000 15.01 9.41 Human
DiDeMo anne2017localizing 1,037 53.94 29.11 Human
MSVD chen2011collecting 670 10.04 7.01 Human
ActivityNet yu2019activitynet 5,044 36.00 13.48 Human
DREAM-1K wang2024tarsier 1,000 8.90 59.30 Human
VDC chai2024auroracap 1,000 28.18 500.91 VLM
CAREBENCH xu2024fine 1,000 14.35 227.95 Human
@@ Abstract
Clip-level 40,804 17.86 1393.65 VLM + Human $\checkmark$
Video-level 467 1560.30 106075.39 VLM + Human $\checkmark$
1 Introduction
In the data-driven era, with the rapid advancement of digital technology, video has become a crucial medium for information transmission due to its intuitive and richly expressive nature pouyanfar2018multimedia . As a core task in multimodal understanding, video-text retrieval, which aims to retrieve semantically relevant video content from large-scale datasets given a textual query, has become increasingly important zhu2023deep .
Recently, videos have grown significantly in length, making retrieving full long-form videos and fine-grained clips increasingly relevant lin2022eclipse ; luo2024video ; eltahir2025multimodal . Consequently, there is an urgent need to benchmark both long video retrieval and clip-level retrieval from long videos to support real-world applications.
However, current video-text retrieval benchmarks are primarily constrained by short video lengths and low-quality captions. These limitations can be summarized in the following three aspects:
@@ 1 Introduction
(1) Limited Video Length Most widely used datasets—such as MSR-VTT xu2016msr , MSVD chen2011collecting , and DiDeMo anne2017localizing contain only short videos, with average durations less than 1 minute. These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
@@
(2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance. While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) malik2021automatic ), these methods often suffer from low accuracy and semantic ambiguity. Although manual annotation wu2024longvideobench ; song2024moviechat ; yu2019activitynet can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets. Balancing scalability with annotation quality remains a critical challenge in current research.
@@
(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos eltahir2025multimodal ; zhu2024video . This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content. In contrast, constructing clip-level benchmarks by segmenting long videos introduces a more realistic and challenging setting—models must not only retrieve relevant content but also distinguish fine-grained clips from densely packed temporal contexts. This setting better reflects practical applications and poses a greater challenge for retrieval models.
@@
To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency. Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) zhang2024vision , (2) automatically assessing caption quality via the EVQAScore model liang2024evqascore , and (3) conducting a final round of human verification. This hybrid process effectively ensures both high fidelity and scalability of annotations. For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model wang2024qwen2 ; bai2025qwen2 , which demonstrates strong performance on multimodal understanding tasks. The integration of EVQAScore liang2024evqascore further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments. By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios. First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content ge2024v2pe ; zhao2023mmicl . As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
Second, captions play a vital role in aligning textual and visual modalities. In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment chen2025expertized . Accurately modeling high-level semantics—such as overarching themes, narrative progression, and emotional tone—remains an open challenge faghihi2024spectrum .
These observations underscore the importance of a dedicated benchmark. By explicitly targeting long-form content, LoVR surfaces technical bottlenecks in current approaches and provides a valuable foundation for developing more robust long video-text retrieval systems.
Our contributions are as follows:
  1. 1.
@@
  3. 3.
    We conduct systematic experiments using various embedding models on LoVR, providing insights into effective representation learning and retrieval strategies for long-form video-text understanding. Our results reveal that LoVR is more challenging than existing benchmarks, especially in the context of long video retrieval, thus introducing a new challenge and setting a fresh research direction for the community.
Table 2: Duration distribution of long videos in LoVR. ‘Avg.’ refers to the average duration in seconds, while ‘#Clip’ indicates the average number of clips obtained by segmenting each video.
Duration (900, 1000] (1000, 2000] (2000, 3000] (3000, 3600]
Statistics Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip
998.7 4 95.5 1293.2 377 88.6 2569.2 61 84.6 3215.6 25 75.1
2 Related Works
@@
Traditional benchmarks (e.g., MSR-VTT xu2016msr , MSVD chen2011collecting , DiDeMo anne2017localizing , TGIF tgif-cvpr2016 , LSMDC lsmdc ) focus on short video clips with concise captions, limiting their ability to evaluate fine-grained semantic understanding. Recent efforts (e.g., ActivityNet yu2019activitynet , HowTo100M miech2019howto100m ) scale up video duration and annotation richness, but HowTo100M relies on noisy ASR-generated captions. To improve annotation quality, DREAM-1K wang2024tarsier and VDC chai2024auroracap use human-annotated descriptions, yet their manual annotation process hinders scalable expansion. To overcome these challenges, we leverage recent advances in multi-concept VLMs an2024mc and visual prompting lin2025drawandunderstand to introduce a VLM-based initial annotator, enabling scalable and fine-grained long-video retrieval.
@@
Recent advances have been driven by deep learning techniques. For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations. zhu2023deeplearningvideotextretrieval ; an2024mc Feature matching has evolved from global(e.g. CLIP4Clip luo2021clip4clipempiricalstudyclip ,VSE++ han2022bicnetlearningefficientspatiotemporal ) to multi-granular approaches (local and individual-level alignment)(e.g. CLIP2TV gao2022clip2tvalignmatchdistill ), improving retrieval accuracy. However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
3 LoVR Datasets Construction
@@
Figure 2: Overview of the three-stage pipeline in LoVR. Step 1 segments long videos into high-dynamic clips using visual change detection and threshold-based filtering. Step 2 generates high-quality clip-level captions via iterative VLM captioning and human fallback based on EVQAScore. Step 3 constructs long-video captions by clustering clip captions and summarizing them into a full-length description. Finally, a human review process is conducted to ensure quality.
3.1 Video Source
The LoVR benchmark is constructed based on the LongVideoBench wu2024longvideobench dataset. LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 3600 seconds (i.e., 1 hour). The domains of the videos can be broadly categorized into four domains: Life, Movie, Knowledge, and News. To focus on long-form video retrieval, we apply a duration filter and retain only those videos exceeding 15 minutes (900 seconds) in length. Table 2 summarizes the distribution of long videos across different duration intervals, including the average duration and the number of videos in each group.
3.2 Dataset Construction Pipeline
@@
3.2.1 Clip Segmentation
Clips serve as fundamental semantic units of videos. Each long video $v\in\mathcal{V}$ is composed of a sequence of temporally contiguous clips, denoted as $v=\{c_{1},c_{2},\dots,c_{N}\}$, where $c_{i}\in\mathcal{C}$. Here, $\mathcal{V}$ denotes the set of all long videos, and $\mathcal{C}$ represents the collection of all clips. Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval. We employ the ContentDetector module from PySceneDetect PySceneDetect , which segments videos based on visual content changes across frames. This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated. We empirically set $\tau=34$, following standard practices in scene detection literature.
Based on visual scene dynamics, videos can be naturally categorized into two types: those with high scene dynamics and those with low scene dynamics. As illustrated in Figure 3, videos such as lectures, interviews, or static photography exhibit minimal visual variation throughout their duration, with scenes remaining largely unchanged across different segments. The inclusion of such videos simplifies the retrieval task, as only a small number of frames may suffice to represent the entire content. In contrast, videos with high scene dynamics contain rich and diverse visual transitions, requiring more precise matching and understanding, which is precisely the level of challenge our benchmark aims to introduce. Furthermore, as shown in the lower portion of Figure 3, the inter-frame difference scores of two representative videos are computed and visualized. The high-scene-dynamics video shows large frame-wise variations, while the low-dynamic one maintains consistently high similarity across frames, often below the threshold for clip segmentation.
@@ 3.2.1 Clip Segmentation
Figure 3: Illustration of Video Dynamics: High vs. Low Scene Dynamics. Example video frames are shown in the upper part (high-dynamic on the left, low-dynamic on the right), while the lower part presents the computed inter-frame difference curves. The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
3.2.2 Caption Generation
@@
We utilize advanced VLMs to produce initial captions based on the visual content of each clip. Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM. To ensure high caption quality, we employ Qwen2.5-VL-Instruct bai2025qwen2 , a series of advanced VLMs. When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style. Here, we emphasize the importance of VLMs in describing theme information, such as overall visual style and atmosphere, which is crucial for understanding videos but often overlooked. As shown in Table 1, existing benchmark captions lack this type of thematic content.
@@
To assess the quality of generated captions, we utilize EVQAScore liang2024evqascore , a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions. For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption. After scoring all instances, we construct a dataset $K=\{(c_{i},s_{i},\text{cap}_{i})\}_{i=1}^{n}$, containing the video clip, its associated caption, and the corresponding quality score.
However, there exists a discrepancy between the computed $s_{i}$ and whether the caption meets human-annotated quality standards. To bridge this gap, we determine an empirical threshold $\theta$ that serves as a benchmark for identifying captions that satisfy manual verification criteria. Specifically, we sample a subset $K^{\prime}\subseteq K$ and conduct manual annotation to count captioning errors per instance. This allows us to establish an empirical correlation between the computed matching scores and human-verified relevance. As illustrated in the left portion of Figure 4, a significant rise in annotation errors is observed for samples with scores below 0.2. Based on this analysis, we set the quality threshold to $\theta=0.2$. Captions with scores above $\theta$ are retained as high-quality, while those below are flagged as low-quality, needing refinement or exclusion.
Figure 4: Left: Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$. The score distribution refers to the left y-axis, while the Content Error Median and Theme Error Median refer to the right y-axis. A significant drop in annotation errors is observed when the match score exceeds 0.2. Middle: Caption length distribution of the long videos. Video-level captions are very long, centered around 10,000 tokens. Right: Caption length distribution of the clips.
@@
For captions identified as low-quality, we switch to an alternative VLM to regenerate captions, which are then re-evaluated using the EVQAScore liang2024evqascore . Only those captions achieving a score above the quality threshold $\theta$ are retained. This strategy leverages the diversity among VLMs—arising from differences in training data and architectural design—to improve caption quality for challenging clips. It is only when the number of failed attempts exceeds three that we resort to manual annotation to generate the final captions. This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
3.2.3 Full Video Captioning

### v2

@@
Abstract.
Long videos contain a vast amount of information, making video-text retrieval an essential and challenging task in multimodal learning and web-scale search. On today’s Web, where users increasingly expect to locate not only relevant pages but also specific long videos or fine-grained clips, existing benchmarks fall short due to limited video duration, low-quality captions, and coarse annotation granularity. To address these limitations, we introduce LoVR, a benchmark specifically designed for long video-text retrieval. LoVR contains 467 long videos and over 40,804 fine-grained clips with high-quality captions. To overcome the issue of poor machine-generated annotations, we propose an efficient caption generation framework that integrates VLM automatic generation, caption quality scoring, and dynamic refinement. This pipeline improves annotation accuracy while maintaining scalability. Furthermore, we introduce a semantic fusion method to generate coherent full-video captions without losing important contextual information. Our benchmark introduces longer videos, more detailed captions, and a larger-scale dataset, presenting new challenges for video understanding and retrieval. Extensive experiments on various advanced models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research. We release the code link at https://anonymous.4open.science/r/LoVR-benchmark-A207/.
Video-Text Retrieval, Multimodal, Benchmark
[…]
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. Example from LoVR. A long video is divided into clips, three shown on the left. The caption on the right describes each clip’s scene and atmosphere, their relations, and the overall summary.
1. Introduction
Table 1. Comparison of statistics of video-text benchmarks.
Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator
MSR-VTT (Xu et al., 2016) 1,000 15.01 9.41 Human
DiDeMo (Anne Hendricks et al., 2017) 1,037 53.94 29.11 Human
MSVD (Chen and Dolan, 2011) 670 10.04 7.01 Human
ActivityNet (Yu et al., 2019) 5,044 36.00 13.48 Human
DREAM-1K (Wang et al., 2024c) 1,000 8.90 59.30 Human
VDC (Chai et al., 2024) 1,000 28.18 500.91 VLM
CAREBENCH (Xu et al., 2024) 1,000 14.35 227.95 Human
@@ 1. Introduction
Clip-level 40,804 17.86 1393.65 VLM + Human
Video-level 467 1560.30 106075.39 VLM + Human
With the explosive growth of online media platforms, video has become a dominant content form on the Web. Notably, long-form videos, such as lectures, documentaries, and live streams lasting from minutes to hours, constitute a significant and growing portion of this information ecosystem. Effective retrieval of such content is therefore essential for modern web search, where users increasingly expect to locate not just relevant videos but specific moments within them. In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (Zhu et al., 2023a).
However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (Lin et al., 2022; Luo et al., 2024; Eltahir et al., 2025). This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(Xu et al., 2016), DiDeMo(Anne Hendricks et al., 2017)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization. Consequently, there is an urgent need for a dedicated benchmark to drive progress in long-form video and clip-level retrieval. The limitations of existing benchmarks can be summarized in three key aspects:
@@
(1) Limited Video Length Most widely used datasets—such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), and DiDeMo (Anne Hendricks et al., 2017) contain only short videos, with average durations less than 1 minute. These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
@@
(2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance. While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (Malik et al., 2021)), these methods often suffer from low accuracy and semantic ambiguity. Although manual annotation (Wu et al., 2024; Song et al., 2024; Yu et al., 2019) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets. Balancing scalability with annotation quality remains a critical challenge in current research.
@@
(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (Eltahir et al., 2025; Zhu et al., 2024). This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content. In contrast, constructing clip-level benchmarks by segmenting long videos introduces a more realistic and challenging setting that models must not only retrieve relevant content but also distinguish fine-grained clips from densely packed temporal contexts. This setting better reflects practical applications and poses a greater challenge for retrieval models.
@@
To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency. Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (Zhang et al., 2024a), (2) automatically assessing caption quality via the EVQAScore (Liang et al., 2024), and (3) conducting a final round of human verification. This hybrid process effectively ensures both high fidelity and scalability of annotations. For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (Wang et al., 2024a; Bai et al., 2025), which demonstrates strong performance on multimodal understanding tasks. The integration of EVQAScore (Liang et al., 2024) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments. By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios. First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (Ge et al., 2024; Zhao et al., 2023). As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser. Second, captions play a vital role in aligning textual and visual modalities. In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (Chen et al., 2025). Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (Faghihi et al., 2024).
These observations underscore the importance of a dedicated benchmark. By explicitly targeting long-form content, LoVR surfaces technical bottlenecks in current approaches and provides a valuable foundation for developing more robust long video-text retrieval systems. Our contributions are as follows:
Figure 2. Overview of the data construction pipeline in LoVR. Step 1 segments long videos into high-dynamic clips using visual change detection and threshold-based filtering. Step 2 generates high-quality clip-level captions via iterative VLM captioning and human fallback based on EVQAScore. Step 3 constructs long-video captions by clustering clip captions and summarizing them into a full-length description. Finally, a human review process is conducted to ensure quality.
  1. (1)
@@
  3. (3)
    We conduct systematic experiments using various models on LoVR, providing insights into effective representation learning and retrieval strategies for long-form video-text understanding. Our results reveal that LoVR is more challenging than existing benchmarks, especially in the context of long video retrieval, thus introducing a new challenge and setting a fresh research direction for the community.
2. Related Works
@@
Traditional benchmarks such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), DiDeMo (Anne Hendricks et al., 2017) TGIF (Li et al., 2016) and LSMDC (Rohrbach et al., 2017) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding. Recent efforts (Yu et al., 2019) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (Miech et al., 2019) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise. To address this limitation, DREAM-1K (Wang et al., 2024c) and VDC (Chai et al., 2024) adopt human-annotated descriptions to improve annotation accuracy. However, their heavy reliance on manual annotation constrains scalable dataset expansion, as the labor-intensive process complicates large-scale data acquisition.
@@
Recent advances have been driven by deep learning techniques. For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations. (Zhu et al., 2023b) Feature matching has evolved from global(e.g. CLIP4Clip (Luo et al., 2021),VSE++ (Han et al., 2022)) to multi-granular approaches (local and individual-level alignment)(e.g. CLIP2TV (Gao et al., 2022)), improving retrieval accuracy. However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
3. LoVR Datasets Construction
@@
Figure 3. Illustration of Video Dynamics: High vs. Low Scene Dynamics. Example video frames are shown in the upper part (high-dynamic on the left, low-dynamic on the right), while the lower part presents the computed inter-frame difference curves. The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
3.1. Video Source
The LoVR benchmark is constructed based on the LongVideoBench (Wu et al., 2024) dataset. LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 1 hour. The domains of the videos can be broadly categorized into four domains: Life, Movie, Knowledge, and News. To focus on long-form video retrieval, we apply a duration filter and retain only those videos exceeding 15 minutes (900 seconds) in length.
3.2. Dataset Construction Pipeline
@@
3.2.1. Clip Segmentation
Clips serve as fundamental semantic units of videos. Each long video $v\in\mathcal{V}$ is composed of a sequence of temporally contiguous clips, denoted as $v=\{c_{1},c_{2},\dots,c_{N}\}$, where $c_{i}\in\mathcal{C}$. Here, $\mathcal{V}$ denotes the set of all long videos, and $\mathcal{C}$ represents the collection of all clips. Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval. We employ the ContentDetector module from PySceneDetect (PySceneDetect, 2021), which segments videos based on visual content changes across frames. This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated.
Based on visual scene dynamics, videos can be naturally categorized into two types: those with high scene dynamics and those with low scene dynamics. As illustrated in Figure 3, videos such as lectures, interviews, or static photography exhibit minimal visual variation throughout their duration, with scenes remaining largely unchanged across different segments. Therefore, we set $\tau=34$, segmenting clips with high-scene dynamics while preserving those with low-scene dynamics.
The inclusion of low-scene dynamic videos simplifies the retrieval task, as only a small number of frames may suffice to represent the entire content. In contrast, videos with high scene dynamics contain rich and diverse visual transitions, requiring more precise matching and understanding, which is precisely the level of challenge our benchmark aims to introduce. Furthermore, as shown in the lower portion of Figure 3, the inter-frame difference scores of two representative videos are computed and visualized. The high-scene-dynamics video shows large frame-wise variations, while the low-dynamic one maintains consistently high similarity across frames, often below the threshold for clip segmentation.
@@ 3.2.1. Clip Segmentation
Figure 4. The figure shows four panels from left to right: (a) Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$. The left y-axis shows the match score distribution, while the right y-axis displays the median content and theme annotation errors. A notable reduction in annotation errors occurs when the match score exceeds 0.2. (b) Distribution of caption lengths for long videos, with most captions concentrated around 10,000 tokens. (c) Distribution of clip durations. (d) Distribution of caption lengths for video clips.
3.2.2. Caption Generation
@@
We utilize advanced VLMs to produce initial captions based on the visual content of each clip. Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM. To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (Bai et al., 2025), a series of advanced VLMs. When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style.
@@
To assess the quality of generated captions, we utilize EVQAScore (Liang et al., 2024), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions. For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption. After scoring all instances, we construct a dataset $K=\{(c_{i},s_{i},\text{cap}_{i})\}_{i=1}^{n}$, containing the video clip, its associated caption, and the corresponding quality score.
However, there exists a discrepancy between the computed $s_{i}$ and whether the caption meets human-annotated quality standards. To bridge this gap, we determine an empirical threshold $\theta$ that serves as a benchmark for identifying captions that satisfy manual verification criteria. Specifically, we sample a subset $K^{\prime}\subseteq K$ and conduct manual annotation to count captioning errors per instance. This allows us to establish an empirical correlation between the computed matching scores and human-verified relevance. As illustrated in Figure 4(a), a significant rise in annotation errors is observed for samples with scores below 0.2. Based on this analysis, we set the quality threshold to $\theta=0.2$. Captions with scores above $\theta$ are retained as high-quality, while those below are flagged as low-quality, needing refinement or exclusion.
@@
For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (Wang et al., 2024a). The initial temperature is set to 0.9 during caption generation. to regenerate captions, which are then re-evaluated using the EVQAScore (Liang et al., 2024). Only those captions achieving a score above the quality threshold $\theta$ are retained. This strategy leverages the diversity among VLMs—arising from differences in training data and architectural design—to improve caption quality for challenging clips. It is only when the number of failed attempts exceeds three that we resort to manual annotation to generate the final captions. This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
3.2.3. Full Video Captioning